please help.
solve the following equation.
1/8 - 1/12 = 1/t That is 1 over 8 minus 1 over 12 equals 1 over t

Answer by priya(51)

try to find the common denominator for the numbers on the left side first..
we get,
(3-2)/24 = 1/t (24 is the common divisor for 8 and 12)
then cross multiply to find t,
we get,
1* t = 24 *1
therefore t = 24
finding difficult to solve pbs..join my online course..visit my site.
priya